OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (1, 3) and is perpendicular to the line y = –x + 2?

OpenStudy (anonymous):

do you know what the slope of \(y=-x+2\) is ?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know what the slope of the perpendicular line would be?

OpenStudy (anonymous):

no, thats what i'm having trouble understanding. Is it -1?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

if the slope of a line is \(m\) then the perpendicular line will have slope \(-\frac{1}{m}\) sometimes called the "negative reciprocal"

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so for example if the slope of a line is 2, the perpendicular line will have slope \(-\frac{1}{2}\)

OpenStudy (anonymous):

and if the slope of a line is \(-\frac{3}{4}\) the perpendicular line will have slope \(\frac{4}{3}\)

OpenStudy (anonymous):

so would it be -1/1?

OpenStudy (anonymous):

your slope is already negative, it is \(-1\) the perpendicular line will have a positive slope

OpenStudy (anonymous):

so 1/1

OpenStudy (anonymous):

yeah, better known as \(1\)

OpenStudy (anonymous):

lol, yeah. ok, i understand, so then how do we get the y-intercept? Do we keep it from the equation? So for instance it would be y=x+2

OpenStudy (anonymous):

?*

OpenStudy (anonymous):

now that you have the slope is 1, and the point is \((1,3)\) use the "point - slope" formula

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] \[y-3=1(x-1)\] \[y=x+2\] yeah you have it it was silly of me to write the 1 there, just to make it clear

OpenStudy (anonymous):

I just did that on my paper! Cool! I fully understand this now, thank you so much for helping me. You're really the best!

OpenStudy (anonymous):

yw